Two electrolytic cells were connected in series.
In one (i) an AgNO3solution and in the other
(ii) 500 cm3of a solution in which 10.0 g of
CuSO4. 5 H2O is dissolved are present. After
passing electric current for 30 minutes, it was
found that 1.307 g silver has been deposited.
Calculate the concentration of copper expressed
in gram per litre (gL-1) in the copper sulphate
solution after the electrolysis.
[Atomic masses are : Cu = 63.54; Ag = 108]
1
Expert's answer
2019-09-17T07:40:48-0400

1)Find the mass of Cu2+ released during electrolysis.

"m_{Cu^{2+}}=\frac{m_{Ag}·M_{Cu}}{M_{Ag}·2}=\frac{1.307·63.54}{108·2}=0.384g"

2) Find the mass of Cu2+

"m_{Cu^{2+}end}=m_{Cu^{2+}}+\frac{m_{CuSO_4·5H_2O}·M_{Cu^{2+}}}{M_{CuSO_4·5H_2O}}=0.384+\frac{10·63.54}{249.69}=2.929g"

3)Find the concentration of Cu2+

"C=\frac{m_{Cu^{2+}end}}{V}=\frac{2.929}{0.5}=5.858g/L"

Answer:5.858g/L